OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):

Standard form equation for a circle with center (h,k) and radius r is: (x−h)2+(y−k)2=r2 So you have the center, now you just need the radius. Well, given the center and a point on the circle, the radius is just the distance between those 2 points. so use the distance formula to find the distance between those 2 points, which is your r: d=(x1−x2)2+(y1−y2)2−−−−−−−−−−−−−−−−−−√ Then plug your center point coordinates, and your r, back into the circle equation that I gave you above.

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2= r^2. d = sqrt [(x1-x2)^2 + (y1-y2)^2]

OpenStudy (anonymous):

SRry forgot to mention square properly